What is the mass of 4.3 moles of water
1
Expert's answer
2016-04-27T10:43:05-0400
n = m/M
where n is amount of moles of water, m is the mass of water, M is a molar mass of water( 18 g/mol).
m = n*M = 4.3*18 = 77.4 g
The mass of 4.3 moles of water is 77.4 g